# Transverse momentum asymmetry in the semi-inclusive electron positron annihilation process

### v1

@@ Abstract
Hadronization, a non-perturbative process, cannot be calculated from first principles. It can be investigated either by phenomenological models or by examining the behavior of produced hadrons or through fragmentation functions. These fragmentation functions are non-perturbative quantities whose determination relies entirely on experimental data. However, higher-twist fragmentation functions present significant challenges for their determination due to power suppression. In this paper, we propose an asymmetry to study twist-3 fragmentation functions. This asymmetry is defined as the transverse momentum asymmetry of the fragmenting quark and/or the produced jet with respect to the observed hadron direction within the semi-inclusive electron positron annihilation process. As a twist-3 effect, this asymmetry is sensitive to the distribution of the jet relative to the produced hadron direction during hadronization. Furthermore, it is closely related to twist-3 transverse momentum dependent fragmentation functions and provides a set of measurable quantities for their determination.
@@ I Introduction
Quantum chromodynamics (QCD) is a fundamental non-Abelian gauge theory of strong interaction whose Lagrangian deals with quarks, gluons, and their interactions. It is characterized by two key features: asymptotic freedom and color confinement. Due to color confinement, isolated quarks and gluons are not experimentally observable; only color-singlet hadrons can be detected. The process in which free quarks and/or gluons combine to form color singlet hadrons is known as hadronization. Hadronization is inherently non-perturbative and cannot be calculated from first principles. Nevertheless, it can be described by phenomenological models, such as Feynman-Field model [1], LUND model [2], Webber model [3, 4], recombination model [5, 6, 7, 8] and so on. While these models provide a global perspective on hadronization, they often lack detailed insights. To address this, one can study the behaviors of the single-hadron or the double-hadron during hadronization process. They respectively correspond to the single-hadron fragmentation functions (FFs) [9, 10], and the double-hadron fragmentation functions (diFFs) [11, 12, 13]. Both FFs and diFFs are non-perturbative quantities; a systematic introduction to them can be found in Ref. [14].
@@
In addition to understanding behaviors of the single-hadron or the double-hadron, it is also very important to study the collective behavior of multiple hadrons during hadronization process. An ensemble of hadrons clustered within a phase space defined by specific parameters constitutes a jet. Jets originating from quarks (or gluons) have received increasing attention in recent years as effective probes of hadronization. Theoretical formulations for jet production in high-energy collisions exhibit simpler forms and avoid introducing additional uncertainties associated with (di)FFs. This is helpful to improve the accuracy of measurements. Theoretical studies of jet production in high-energy collisions encompass two approaches: one focuses on higher-twist calculations up to leading order [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], while the other focuses on higher-order calculations [26, 27, 31, 30, 28, 29, 33, 32]. Furthermore, the introduction of jet fragmentation functions (JFFs) [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], designed to study hadron distributions within a jet, has significantly broadened the research area. An observed hadron carries not only a longitudinal momentum fraction relative to the jet but also a transverse momentum relative to the jet axis. Consequently, JFFs encode a set of angular modulations of the hadron distributions within the jet, providing valuable insights into the hadronization process. Conversely, in the hadron collinear frame, the jet axis acquires transverse momentum ($\vec{j}_{T}$) relative to the direction of the observed hadron. Under the approximation that the jet axis aligns with the direction of the initial parton (quark or antiquark), the intrinsic transverse momentum of the parton satisfies $\vec{k}^{\prime}_{T}=\vec{j}_{T}$. Therefore, unlike jet fragmentation functions (JFFs), which probe the distributions of a hadron within the jet, $\vec{j}_{T}$ serves as an effective probe for studying jet distributions relative to the observed hadron direction.
@@
To be explicit, we organize this paper as follows. In Sec. II, we present a introduction of the reference frame used in this paper and show the definition of the transverse momentum asymmetry in this frame. In Sec. III, we present the general theoretical framework for calculating the hadronic tensor and the differential cross section up to twist-3. In Sec. IV, we perform the numerical estimates for the transverse momentum asymmetry. Finally, we give a brief summary in Sec. V.
@@ II Transverse momentum asymmetry
where the momenta of the relevant particles are given in parentheses. To define the transverse momentum asymmetry, we need to establish a reference frame and determine the transverse direction. We begin by introducing the center-of-mass frame of the leptons, where the observed hadron travels along the +z+z direction and the momenta of leptons lie in the xzx-z plane, also referred to as the reaction plane (see Fig. 1). The transverse direction is then defined as the direction perpendicular to the $z$ axis, lying within the xyx-y plane. In this frame, we can introduce the light-cone unit vectors, $\bar{n}^{\mu}=(1,0,\vec{0}_{T})$ and $n^{\mu}=(0,1,\vec{0}_{T})$, which satisfy $\bar{n}^{2}=n^{2}=0$ and $\bar{n}\cdot n=1$. These vectors allow us to define the transverse tensors for further analysis,
@@
To decompose the leptonic tensor illustrated in the following section, we also introduce the time-like vector $\hat{t}^{\mu}$ and the space-like vector $\hat{v}^{\mu}$,
@@
In the center-of-mass frame of the leptons illustrated in Fig. 1, quark and anti-quark are produced back-to-back at tree level without higher-order gluon radiations (two-jet event). Their momenta are labeled as $k$ and $k^{\prime}$, respectively. In the jet production semi-inclusive annihilation considered in this paper, we take $\vec{k}^{\prime}$ as the momentum of the jet under the approximation that the jet axis direction coincides with that of the antiquark. The hadron is detected in the back-to-back hemisphere. Given that the transverse momentum is defined in the xyx-y plane, the transverse momentum asymmetry can be expressed as a difference of the jet transverse momentum ($k_{\perp}^{\prime}$). Therefore, we have
@@ III Theoretical framework
Here, $e_{l}$ and $e_{q}$ are electric charges of relevant leptons and quarks, $M_{Z}$ and $\Gamma_{Z}$ are respectively the mass and width of the $Z^{0}$ boson, $\theta_{W}$ is the weak mixing angle.
The leptonic tensors for the electromagnetic term, interference term, and the weak term are given by
@@
where $\lambda_{e}$ is the helicity of the initial electron, and $c_{V}^{e}$ and $c_{A}^{e}$ are defined in the weak current $J^{\mu}=\bar{\psi}\gamma^{\mu}(c_{V}^{e}-c_{A}^{e}\gamma^{5})\psi$ with $c_{1}^{e}=(c_{V}^{e})^{2}+(c_{A}^{e})^{2}$ and $c_{3}^{e}=2c_{v}^{e}c_{A}^{e}$. We can rewrite the leptonic tensor in terms of vectors $\hat{t}$ and $\hat{v}$ as follows,
@@
where unit vector $\hat{x}^{\mu}$ is defined as $\hat{x}^{\mu}=l_{T}^{\mu}/QD(y)$. For simplicity, the functions of $y$ are written as
@@
The expression for the leptonic tensor in Eq. (25) can greatly simplify the contraction with the hadronic tensor in the framework of parton model.
@@ III.1 Hadronic tensor in the parton model
Hadronic tensor contains non-perturbative information and can be calculated in terms of the structure functions [18, 46]. Given that our focus is on transverse momentum effects and twist-3 FFs, we do not introduce structure functions and instead restrict the analysis to calculation within the parton model framework.
@@ III.2 Fragmentation functions
Correlators shown in Eqs. (37) and (38) are $4\times 4$ matrices that can be decomposed by Dirac matrices. To satisfy the helicity conservation, we only need to consider the chiral-even terms, $\gamma^{\alpha}$- and the $\gamma^{5}\gamma^{\alpha}$-terms, when decomposing the correlators. For example, $\hat{\Xi}^{(0)}=\gamma^{\alpha}\Xi_{\alpha}^{(0)}+\gamma^{5}\gamma^{\alpha}\tilde{\Xi}_{\alpha}^{(0)}$. Here $\Xi_{\alpha}^{(0)}$ and $\tilde{\Xi}_{\alpha}^{(0)}$ are coefficient functions, which can be obtained by
@@ III.3 Calculations up to twist-3
The contribution of the leading twist comes only from the quark-quark correlator $\hat{\Xi}^{(0)}$. To calculate it, we use Eqs. (41) and (42), and the following relations
@@
Nevertheless, twist-3 contributions have two sources, one is the quark-quark correlator $\hat{\Xi}^{(0)}$ and the other is the quark-gluon-quark correlator $\hat{\Xi}^{(1)}_{\rho}$. We first calculate these contributions from the quark-quark correlator $\hat{\Xi}^{(0)}$. Here we use the following traces
@@
For the twist-3 contribution from the correlator $\hat{\Xi}^{(1)}_{\rho}$, we use the following traces,
@@
and obtain the following result,
@@
We have provided the detailed calculation for obtaining this expression in the appendix A.
@@ III.4 Differential cross section
Other contractions for twist-3 terms can be obtained similarly.
@@
Up to now, we have presented the results of the differential cross section using the weak interaction term as an example. To obtain the results for electromagnetic term, we need to set $c_{3}^{e/q}=0$ and $c_{1}^{e/q}=1$. For the interference term, we need to set $c_{3}^{e/q}=c_{A}^{e/q}$ and $c_{1}^{e/q}=c_{V}^{e/q}$. The kinematic factors are also different. To make it transparent, we can get the electromagnetic and interference cross sections by replacing the parameters in the weak interaction cross section according to Tab. 1
@@
Table 1: Relations of kinematic factors between weak, EM, and interference interactions.
@@ IV Numerical estimates of the transverse momentum asymmetry
As introduced in Sec. II, the transverse momentum asymmetry is defined through the difference of the transverse momenta, see Eqs. (16) and (17). Note that the transverse momentum of the fragmenting quark lies in the xyx-y plane, it can therefore be decomposed as
@@
Therefore, based on the expressions of the differential cross sections in Eq. (60) and Eq. (65), we can introduce the general definitions of the transverse momentum asymmetries,
$A^{x}$ $=\frac{\int_{-\pi/2}^{\pi/2}d\phi~d\tilde{\sigma}-\int_{\pi/2}^{3\pi/2}d\phi~d\tilde{\sigma}}{\int_{-\pi/2}^{\pi/2}d\phi~d\tilde{\sigma}_{U}+\int_{\pi/2}^{3\pi/2}d\phi~d\tilde{\sigma}_{U}},$ (74)
$A^{y}$ $=\frac{\int_{0}^{\pi}d\phi~d\tilde{\sigma}-\int_{\pi}^{2\pi}d\phi~d\tilde{\sigma}}{\int_{0}^{\pi}d\phi~d\tilde{\sigma}_{U}+\int_{\pi}^{2\pi}d\phi~d\tilde{\sigma}_{U}}.$ (75)
Here $A^{x}$ and $A^{y}$ can lead to asymmetries in the $x$-direction and $y$-direction, respectively.
According to the definitions above, four kinds of asymmetries are obtained. They are
$A_{U}^{x}=-\frac{8\kappa_{M}k_{\perp M}}{z\pi}\frac{A_{ZZ}}{A_{ZZ}}\frac{\left[T^{q}_{2}(y)-\lambda_{e}\tilde{T}^{q}_{2}(y)\right]}{\left[T^{q}_{0}(y)-\lambda_{e}\tilde{T}^{q}_{0}(y)\right]}\frac{D^{\perp}}{D_{1}},$ (76)
$A_{U}^{y}=+\frac{8\kappa_{M}k_{\perp M}}{z\pi}\frac{A_{ZZ}}{A_{ZZ}}\frac{\left[T^{q}_{3}(y)-\lambda_{e}\tilde{T}^{q}_{3}(y)\right]}{\left[T^{q}_{0}(y)-\lambda_{e}\tilde{T}^{q}_{0}(y)\right]}\frac{G^{\perp}}{D_{1}},$ (77)
$A_{L}^{x}=-\frac{8\kappa_{M}k_{\perp M}}{z\pi}\frac{A_{ZZ}}{A_{ZZ}}\frac{\left[T^{q}_{3}(y)-\lambda_{e}\tilde{T}^{q}_{3}(y)\right]}{\left[T^{q}_{0}(y)-\lambda_{e}\tilde{T}^{q}_{0}(y)\right]}\frac{G_{L}^{\perp}}{D_{1}},$ (78)
$A_{L}^{y}=-\frac{8\kappa_{M}k_{\perp M}}{z\pi}\frac{A_{ZZ}}{A_{ZZ}}\frac{\left[T^{q}_{2}(y)-\lambda_{e}\tilde{T}^{q}_{2}(y)\right]}{\left[T^{q}_{0}(y)-\lambda_{e}\tilde{T}^{q}_{0}(y)\right]}\frac{D_{L}^{\perp}}{D_{1}},$ (79)
where the subscripts $U$ and $L$ denote unpolarized and the longitudinally polarized hadrons, respectively. We note that these transverse momentum asymmetries are all twist-3 measurable quantities. We note again that only weak interaction results are shown in Eqs. (76)-(79). For the complete results, electromagnetic and interference terms should be included.
@@
$A_{U}^{x}=-\frac{8\kappa_{M}k_{\perp M}}{z\pi}\frac{e_{q}^{2}}{e_{q}^{2}}\frac{C(y)D(y)}{A(y)}\frac{D^{\perp}}{D_{1}},$ (80)
$A_{U}^{y}=-\frac{8\kappa_{M}k_{\perp M}}{z\pi}\frac{e_{q}^{2}}{e_{q}^{2}}\frac{\lambda_{e}D(y)}{A(y)}\frac{G^{\perp}}{D_{1}},$ (81)
$A_{L}^{x}=+\frac{8\kappa_{M}k_{\perp M}}{z\pi}\frac{e_{q}^{2}}{e_{q}^{2}}\frac{\lambda_{e}D(y)}{A(y)}\frac{G_{L}^{\perp}}{D_{1}},$ (82)
$A_{L}^{y}=-\frac{8\kappa_{M}k_{\perp M}}{z\pi}\frac{e_{q}^{2}}{e_{q}^{2}}\frac{C(y)D(y)}{A(y)}\frac{D_{L}^{\perp}}{D_{1}}.$ (83)
@@
  • Figure 5 shows that asymmetry $A_{U}^{x}$ is not sensitive to fraction $z$. From Eqs. (16) and (17), we notice that the asymmetry is defined through the difference of the transverse momenta, and it should therefore be independent of the longitudinal momentum fraction $z$. Certainly, this conclusion strongly depends on the parameterized form of $D^{\perp}$ in Eq. (85), which imply that $D^{\perp}$ maybe proportional to $zD_{1}$.
@@
We also check that asymmetry $A_{U}^{x}$ is not sensitive to the mean squared transverse momenta of TMD FFs. Furthermore, the use of NNFF10 [56] for $\pi^{+}$, MAPFF10 [57] for $\pi^{+}$, and NPC23 [47, 48] for $K^{+}$ yields consistent results. These findings suggest that the transverse momentum asymmetry is very likely a universal physical quantity that appear to be independent of the identity of the observed hadronic state. It reveals the transverse momentum effects and does not depend on the longitudinal momentum fraction.

### v2

@@ Abstract
Hadronization, a nonperturbative process, cannot be calculated from first principles. It can be investigated either by using phenomenological models or by examining the behavior of produced hadrons or through fragmentation functions. These fragmentation functions are nonperturbative quantities whose determination relies entirely on experimental data. However, higher-twist fragmentation functions present significant challenges for their determination due to power suppression. In this paper, we propose an asymmetry to study twist-3 fragmentation functions. This asymmetry is defined as the transverse momentum asymmetry of the fragmenting quark and/or the produced jet with respect to the observed hadron direction within the semi-inclusive electron positron annihilation process. As a twist-3 effect, this asymmetry is sensitive to the distribution of the jet relative to the produced hadron direction during hadronization. Furthermore, it is closely related to twist-3 transverse momentum dependent fragmentation functions and provides a set of measurable quantities for their determination.
@@ I Introduction
Quantum chromodynamics (QCD) is a fundamental non-Abelian gauge theory of strong interaction whose Lagrangian deals with quarks, gluons, and their interactions. It is characterized by two key features: asymptotic freedom and color confinement. Due to color confinement, isolated quarks and gluons are not experimentally observable; only color-singlet hadrons can be detected. The process in which free quarks and/or gluons combine to form color singlet hadrons is known as hadronization. Hadronization is inherently nonperturbative and cannot be calculated from first principles. Nevertheless, it can be described by phenomenological models, such as the Feynman-Field model [1], the LUND model [2], the Webber model [3, 4], and the recombination model [5, 6, 7, 8]. While these models provide a global perspective on hadronization, they often lack detailed insights. To address this, one can study the behaviors of the single hadron or the double-hadron during hadronization process. They respectively correspond to the single-hadron fragmentation functions (FFs) [9, 10] and the double-hadron fragmentation functions (diFFs) [11, 12, 13]. Both FFs and diFFs are nonperturbative quantities; a systematic introduction to them can be found in Ref. [14].
@@
In addition to understanding behaviors of the single hadron or the double hadron, it is also very important to study the collective behavior of multiple hadrons during the hadronization process. An ensemble of hadrons clustered within a phase space defined by specific parameters constitutes a jet. Jets originating from quarks (or gluons) have received increasing attention in recent years as effective probes of hadronization. Theoretical formulations for jet production in high-energy collisions exhibit simpler forms and avoid introducing additional uncertainties associated with (di)FFs. This is helpful to improve the accuracy of measurements. Theoretical studies of jet production in high-energy collisions encompass two approaches: one focuses on higher-twist calculations up to leading order [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], while the other focuses on higher-order calculations [26, 27, 31, 30, 28, 29, 33, 32]. Furthermore, the introduction of jet fragmentation functions (JFFs) [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], designed to study hadron distributions within a jet, has significantly broadened the research area. An observed hadron carries not only a longitudinal momentum fraction relative to the jet but also a transverse momentum relative to the jet axis. Consequently, JFFs encode a set of angular modulations of the hadron distributions within the jet, providing valuable insights into the hadronization process. Conversely, in the hadron collinear frame, the jet axis acquires transverse momentum ($\vec{j}_{T}$) relative to the direction of the observed hadron. Under the approximation that the jet axis aligns with the direction of the initial parton (quark or antiquark), the intrinsic transverse momentum of the parton satisfies $\vec{k}^{\prime}_{T}=\vec{j}_{T}$. Therefore, unlike JFFs, which probe the distributions of a hadron within the jet, $\vec{j}_{T}$ serves as an effective probe for studying jet distributions relative to the observed hadron direction.
@@
To be explicit, we organize this paper as follows. In Sec. II, we present an introduction of the reference frame used in this paper and show the definition of the transverse momentum asymmetry in this frame. In Sec. III, we present the general theoretical framework for calculating the hadronic tensor and the differential cross section up to twist-3. In Sec. IV, we perform the numerical estimates for the transverse momentum asymmetry. Finally, we give a brief summary in Sec. V.
@@ II Transverse momentum asymmetry
where the momenta of the relevant particles are given in parentheses. To define the transverse momentum asymmetry, we need to establish a reference frame and determine the transverse direction. We begin by introducing the center-of-mass frame of the leptons, where the observed hadron travels along the +z+z direction and the momenta of leptons lie in the $x$-$z$ plane, also referred to as the reaction plane (see Fig. 1). The transverse direction is then defined as the direction perpendicular to the $z$ axis, lying within the $x$-$y$ plane. In this frame, we can introduce the light-cone unit vectors, $\bar{n}^{\mu}=(1,0,\vec{0}_{T})$ and $n^{\mu}=(0,1,\vec{0}_{T})$, which satisfy $\bar{n}^{2}=n^{2}=0$ and $\bar{n}\cdot n=1$. These vectors allow us to define the transverse tensors for further analysis,
@@
To decompose the leptonic tensor illustrated in the following section, we also introduce the timelike vector $\hat{t}^{\mu}$ and the spacelike vector $\hat{v}^{\mu}$,
@@
In the center-of-mass frame of the leptons illustrated in Fig. 1, the quark and anti-quark are produced back-to-back at tree level without higher-order gluon radiations (two-jet event). Their momenta are labeled as $k$ and $k^{\prime}$, respectively. In the jet production semi-inclusive annihilation considered in this paper, we take $\vec{k}^{\prime}$ as the momentum of the jet under the approximation that the jet axis direction coincides with that of the antiquark. The hadron is detected in the back-to-back hemisphere. Given that the transverse momentum is defined in the $x$-$y$ plane, the transverse momentum asymmetry can be expressed as a difference of the jet transverse momentum ($k_{\perp}^{\prime}$). Therefore, we have
@@ III Theoretical framework
Here, $e_{l}$ and $e_{q}$ are electric charges of relevant leptons and quarks, $M_{Z}$ and $\Gamma_{Z}$ are, respectively, the mass and width of the $Z^{0}$ boson, and $\theta_{W}$ is the weak mixing angle.
The leptonic tensors for the electromagnetic term, the interference term, and the weak term are given by
@@
where $\lambda_{e}$ is the helicity of the initial electron, and $c_{V}^{e}$ and $c_{A}^{e}$ are defined in the weak current $J^{\mu}=\bar{\psi}\gamma^{\mu}(c_{V}^{e}-c_{A}^{e}\gamma^{5})\psi$ with $c_{1}^{e}=(c_{V}^{e})^{2}+(c_{A}^{e})^{2}$ and $c_{3}^{e}=2c_{v}^{e}c_{A}^{e}$. We can rewrite the leptonic tensor in terms of vectors $\hat{t}$ and $\hat{v}$ as follows:
@@
where the unit vector $\hat{x}^{\mu}$ is defined as $\hat{x}^{\mu}=l_{T}^{\mu}/QD(y)$. For simplicity, the functions of $y$ are written as
@@
The expression for the leptonic tensor in Eq. (25) can greatly simplify the contraction with the hadronic tensor in the framework of the parton model.
@@ III.1 Hadronic tensor in the parton model
The hadronic tensor contains nonperturbative information and can be calculated in terms of the structure functions [18, 46]. Given that our focus is on transverse momentum effects and twist-3 FFs, we do not introduce structure functions and instead restrict the analysis to calculation within the parton model framework.
@@ III.2 Fragmentation functions
Correlators shown in Eqs. (37) and (38) are $4\times 4$ matrices that can be decomposed by Dirac matrices. To satisfy the helicity conservation, we only need to consider the chiral-even terms, $\gamma^{\alpha}$, and the $\gamma^{5}\gamma^{\alpha}$ terms, when decomposing the correlators. For example, $\hat{\Xi}^{(0)}=\gamma^{\alpha}\Xi_{\alpha}^{(0)}+\gamma^{5}\gamma^{\alpha}\tilde{\Xi}_{\alpha}^{(0)}$. Here $\Xi_{\alpha}^{(0)}$ and $\tilde{\Xi}_{\alpha}^{(0)}$ are coefficient functions, which can be obtained by
@@ III.3 Calculations up to twist-3
The contribution of the leading twist comes only from the quark-quark correlator $\hat{\Xi}^{(0)}$. To calculate it, we use Eqs. (41) and (42), and the following relations:
@@
Nevertheless, twist-3 contributions have two sources, one is the quark-quark correlator $\hat{\Xi}^{(0)}$ and the other is the quark-gluon-quark correlator $\hat{\Xi}^{(1)}_{\rho}$. We first calculate these contributions from the quark-quark correlator $\hat{\Xi}^{(0)}$. Here we use the following traces:
@@
For the twist-3 contribution from the correlator $\hat{\Xi}^{(1)}_{\rho}$, we use the following traces:
@@
and obtain the following result:
@@
We have provided the detailed calculation for obtaining this expression in Appendix A.
@@ III.4 Differential cross section
Other contractions for twist-3 terms can be obtained similarly. We do not show them here for simplicity.
@@
Up to now, we have presented the results of the differential cross section using the weak interaction term as an example. To obtain the results for the electromagnetic term, we need to set $c_{3}^{e/q}=0$ and $c_{1}^{e/q}=1$. For the interference term, we need to set $c_{3}^{e/q}=c_{A}^{e/q}$ and $c_{1}^{e/q}=c_{V}^{e/q}$. The kinematic factors are also different. To make it transparent, we can get the electromagnetic and interference cross sections by replacing the parameters in the weak interaction cross section according to Table. 1
@@
Table 1: Relations of kinematic factors between weak, electromagnetic, and interference interactions.
@@ IV Numerical estimates of the transverse momentum asymmetry
As introduced in Sec. II, the transverse momentum asymmetry is defined through the difference of the transverse momenta, see Eqs. (16) and (17). Note that the transverse momentum of the fragmenting quark lies in the $x$-$y$ plane, and it can be decomposed as
@@
Therefore, $k_{\perp}^{x}(x>0)$ and $k_{\perp}^{x}(<0)$ can be related to the differential cross section via the following equations:
$k_{\perp}^{x}(x>0)$ $\to\int_{x>0}\vec{k}_{\perp}d^{2}k_{\perp}\frac{d\sigma}{d^{2}k_{\perp}}=\int_{x>0}k_{\perp}^{2}\cos\phi dk_{\perp}d\phi\frac{d\sigma}{d^{2}k_{\perp}}$
$=\int^{\pi/2}_{-\pi/2}k_{\perp}^{2}\cos\phi dk_{\perp}d\phi\frac{d\sigma}{d^{2}k_{\perp}dzdy}dzdy,$ (74)
$k_{\perp}^{x}(x>0)$ $\to\int_{x>0}\vec{k}_{\perp}d^{2}k_{\perp}\frac{d\sigma}{d^{2}k_{\perp}}=\int_{x>0}k_{\perp}^{2}\cos\phi dk_{\perp}d\phi\frac{d\sigma}{d^{2}k_{\perp}}$
$=\int^{3\pi/2}_{\pi/2}k_{\perp}^{2}\cos\phi dk_{\perp}d\phi\frac{d\sigma}{d^{2}k_{\perp}dzdy}dzdy,$ (75)
where $\frac{d\sigma}{d^{2}k_{\perp}dzdy}$ is the complete differential cross section shown in Eqs. (60) and (65). For the production of the unpolarized hadron, we therefore have
$k_{\perp}^{x}(x>0)$ $=\frac{2\alpha A_{ZZ}}{Q^{2}}\left[T^{q}_{0}(y)-\lambda_{e}\tilde{T}^{q}_{0}(y)\right]k_{\perp}^{2}D_{1}$
$-\frac{\alpha A_{ZZ}}{Q^{2}}\frac{2}{zQ}\left[T^{q}_{2}(y)-\lambda_{e}\tilde{T}^{q}_{2}(y)\right]k_{\perp}^{3}D^{\perp},$ (76)
$k_{\perp}^{x}(x<0)$ $=\frac{2\alpha A_{ZZ}}{Q^{2}}\left[T^{q}_{0}(y)-\lambda_{e}\tilde{T}^{q}_{0}(y)\right]k_{\perp}^{2}D_{1}$
$+\frac{\alpha A_{ZZ}}{Q^{2}}\frac{2}{zQ}\left[T^{q}_{2}(y)-\lambda_{e}\tilde{T}^{q}_{2}(y)\right]k_{\perp}^{3}D^{\perp}.$ (77)
According to the definitions shown in Eqs. (16) and (17) and the results shown above, four kinds of asymmetries are expressed as
$A_{U}^{x}=-\frac{\kappa_{M}k_{\perp M}}{z}\frac{A_{ZZ}}{A_{ZZ}}\frac{\left[T^{q}_{2}(y)-\lambda_{e}\tilde{T}^{q}_{2}(y)\right]}{\left[T^{q}_{0}(y)-\lambda_{e}\tilde{T}^{q}_{0}(y)\right]}\frac{D^{\perp}}{D_{1}},$ (78)
$A_{U}^{y}=+\frac{\kappa_{M}k_{\perp M}}{z}\frac{A_{ZZ}}{A_{ZZ}}\frac{\left[T^{q}_{3}(y)-\lambda_{e}\tilde{T}^{q}_{3}(y)\right]}{\left[T^{q}_{0}(y)-\lambda_{e}\tilde{T}^{q}_{0}(y)\right]}\frac{G^{\perp}}{D_{1}},$ (79)
$A_{L}^{x}=-\frac{\kappa_{M}k_{\perp M}}{z}\frac{A_{ZZ}}{A_{ZZ}}\frac{\left[T^{q}_{3}(y)-\lambda_{e}\tilde{T}^{q}_{3}(y)\right]}{\left[T^{q}_{0}(y)-\lambda_{e}\tilde{T}^{q}_{0}(y)\right]}\frac{G_{L}^{\perp}}{D_{1}},$ (80)
$A_{L}^{y}=-\frac{\kappa_{M}k_{\perp M}}{z}\frac{A_{ZZ}}{A_{ZZ}}\frac{\left[T^{q}_{2}(y)-\lambda_{e}\tilde{T}^{q}_{2}(y)\right]}{\left[T^{q}_{0}(y)-\lambda_{e}\tilde{T}^{q}_{0}(y)\right]}\frac{D_{L}^{\perp}}{D_{1}},$ (81)
where the subscripts $U$ and $L$ denote unpolarized and the longitudinally polarized hadrons, respectively. We note that these transverse momentum asymmetries are all twist-3 measurable quantities. We note again that only weak interaction results are shown in Eqs. (78)–(81). For the complete results, electromagnetic and interference terms should be included.
@@
$A_{U}^{x}=-\frac{\kappa_{M}k_{\perp M}}{z}\frac{e_{q}^{2}}{e_{q}^{2}}\frac{C(y)D(y)}{A(y)}\frac{D^{\perp}}{D_{1}},$ (82)
$A_{U}^{y}=-\frac{\kappa_{M}k_{\perp M}}{z}\frac{e_{q}^{2}}{e_{q}^{2}}\frac{\lambda_{e}D(y)}{A(y)}\frac{G^{\perp}}{D_{1}},$ (83)
$A_{L}^{x}=+\frac{\kappa_{M}k_{\perp M}}{z}\frac{e_{q}^{2}}{e_{q}^{2}}\frac{\lambda_{e}D(y)}{A(y)}\frac{G_{L}^{\perp}}{D_{1}},$ (84)
$A_{L}^{y}=-\frac{\kappa_{M}k_{\perp M}}{z}\frac{e_{q}^{2}}{e_{q}^{2}}\frac{C(y)D(y)}{A(y)}\frac{D_{L}^{\perp}}{D_{1}}.$ (85)
@@
  • Figure 5 shows that asymmetry $A_{U}^{x}$ is not sensitive to fraction $z$. From Eqs. (16) and (17), we notice that the asymmetry is defined through the difference of the transverse momenta, and it should therefore be independent of the longitudinal momentum fraction $z$. Certainly, this conclusion strongly depends on the parametrized form of $D^{\perp}$ in Eq. (87), which implies that $D^{\perp}$ may be proportional to $zD_{1}$.
We also check that asymmetry $A_{U}^{x}$ is not sensitive to the mean squared transverse momenta of TMD FFs. Furthermore, the use of NNFF10 [56] for $\pi^{+}$, MAPFF10 [57] for $\pi^{+}$, and NPC23 [47, 48] for $K^{+}$ yields consistent results. These findings suggest that the transverse momentum asymmetry is very likely a universal physical quantity that appears to be independent of the identity of the observed hadronic state. It reveals the transverse momentum effects and does not depend on the longitudinal momentum fraction.